OpenStudy (s117):

Which pair of ratios is proportional? A. 4/9 and 16/45 B. 5/12 and 25/144 C. 12/18 and 18/27 D. 15/20 and 17/10

OpenStudy (kittiwitti1):

proportion and cross multiply

OpenStudy (igreen):

You can cross multiply and see if they equal each other.. \(\dfrac{4}{9} = \dfrac{16}{45}\) \(\dfrac{4}{9} \nearrow \dfrac{16}{45}\) \(\dfrac{4}{9} \searrow \dfrac{16}{45}\)

OpenStudy (kittiwitti1):

\[\text{first one: }[4\times 45=9\times 16]?\]

OpenStudy (s117):

first one = 144/180

OpenStudy (kittiwitti1):

no

OpenStudy (kittiwitti1):

does 4 x 45 = 9 x 16

OpenStudy (s117):

fliped it srry 180/144

OpenStudy (kittiwitti1):

._.

OpenStudy (igreen):

That gives you 180 = 144, not 180/144 And 180 doesn't equal 144, so the first one is incorrect.

OpenStudy (kittiwitti1):

\[180\ne 144\]

OpenStudy (igreen):

Try: \(\dfrac{12}{18} = \dfrac{18}{27}\)

OpenStudy (s117):

k

OpenStudy (kittiwitti1):

\[\text{does }12\times 27=18\times 18?\]

OpenStudy (s117):

324 for both

OpenStudy (kittiwitti1):

Then you have your answer :)

OpenStudy (s117):

thnx @iGreen and @kittiwitti1

OpenStudy (igreen):

Yep, so C's our answer.

OpenStudy (kittiwitti1):

When you put the fractions in a proportion and the cross-multiplication values are equal, you have two equivalent fractions. You're welcome :)